# Objective-Free Local Learning and Emergent Language Structure in Thinking Machines

### v1

@@ Abstract.
We introduce a neuro-symbolic framework for generative language modeling, based on local event-drive emergent learning. The framework is centered around hierarchical Hopfield memory chains, which act as a compositional short-term memory & dynamic tokenizer (def. retokenizer). Rather than relying on predefined tokens or supervised training, the model builds structure from scratch, learning to retokenize symbol sequences into multi-scale, compositional representations. It does this by hierarchically constructing projection tensors, which glue together symbol sequences into multi-scale tokens. This imbues the model with an inherit redundancy (i.e an emergent gauge structure), and allows the model to compress local feature co-activations into long-range dependencies.
Curiously, we find that the retokenizer can filter natural language patterns from noise, generating synthetic languages with coherent internal morphology – quantifiably the same as human language. Language is learned in a local (Hebbian) fashion, where model constraints dictate allowed emergent structure, and new information is retained in alignment with this structure.
The lack of global objective allows for a type of plasticity not seen in other language models, including allowing it to generalize beyond its initial inference class (and can do so without explicit data). We show that briefly exposing a single new neuron to the retokenizer during inference, binds the distributed multi-scale token features into a symbolic embedding neuron. These embedding neurons behave as an emergent long-term memory, and introduces a key-value memory mechanism that enables compositional inference and generalization.
This architecture provides a methodological foundation for studying how symbolic structure can emerge from local neural learning. It offers a new pathway for building scalable, interpretable neuro-symbolic systems — where tokens, grammar, and reasoning arise as compressed memory traces within a fully unsupervised Hopfield hierarchy. This approach advances the development of scalable, neuromorphic architectures for generative language models.
Exotic AI, Hebbian, Hopfield Networks, Physics-inspired Neural Nets, Interpretable Models, Hierarchical Compression, Language Emergence, Neuro-symbolic, Neuromorphic
@@ 1. Introduction
Objective-free local learning is the paradigm by which much of the observed structure and organization in the universe arises. In this setting, learning is governed solely by local rules—no global objective functions, no central coordination. Such systems evolve asynchronously, with higher-order and longer-range correlations emerging as incidental consequences of local interactions unfolding over time.
This paradigm includes many natural and artificial systems: human cognition, biochemical self-organization, and even simple cellular automata such as Conway’s Game of Life. Crucially, learning can emerge from random initial conditions without prestructured input. This implies that structure can be learned in a truly data-free regime, where the learning system constructs its own vocabulary and concepts directly from noise.
In a companion paper (Eugenio, 2025), we derived an effective human language model based on this principle: a hierarchical-chain Hopfield network trained via local Hebbian updates. There, we showed how unsupervised hierarchical feature construction can filter linguistic patterns from noise, giving rise to morphologically coherent vocabularies without supervision or data.
Here, we lay out the broader methodological framework inspired by that model. We describe how local learning events can organize into a recursive projection hierarchy, how memory emerges through replay, and how the resulting system supports key capacities of generative language modeling—tokenization, compositional structure, fast autoregressive inference, and compression—without requiring a global objective, backpropagation, nor explicit data.
This framework also bridges two traditionally distinct approaches to language modeling: correlation-based neural systems (Vaswani et al., 2017; Kaplan et al., 2020; Ramsauer et al., 2021; Kozachkov et al., 2023; Hopfield, 1982; Schultz Kisby et al., 2024) and symbolic compositional reasoning (Lake et al., 2017; d’Avila Garcez and Lamb, 2020; Colelough and Regli, 2025; Lamb et al., 2021; West et al., 2024; Dong et al., 2019). At the lower levels, the model constructs a directed acyclic graph (DAG) of valid $n$-gram combinations, gradually building a recursively compositional vocabulary through a process we call retokenization. Each layer of the hierarchy learns local statistical correlations–akin to identifying relevant interactions in a renormalization group flow (Wilson and Kogut, 1974; Shankar, 1994, 2017)–and projects them into higher-level tokens. In this way, the system builds structured representations across scales, where short-range associations give rise to long-range compositional units.
This sub-symbolic mechanism allows the model to infer token boundaries without explicit markers (e.g space or stop tokens), defining a notion of string acceptance grounded in learned statistical compatibility. The projector maps encode a hierarchy of memory that approximates automaton-like behavior [^1] Formally, regular languages are recognized by finite-state automata that lack explicit memory for counting or nesting (Hopcroft et al., 2001). The early stages of our model resemble finite-state structures through local $n$-gram interactions. However, continual learning causes the model to generalize beyond its initial inference class, and may implicitly support more complex behaviors beyond finite-state expressivity (Pérez et al., 2019; Qiu et al., 2025; Rendell, 2011; Pérez et al., 2021; Rendell, 2014; Hahn, 2020; Weiss et al., 2018). A rigorous classification of the exact formal grammar class supported by this model remains a promising direction for future work.: a sequence is tokenized only if it conforms to the model’s internal structure, and next-token predictions follow trajectories constrained by these memory-induced transitions. Thus, inference behaves like a grammar operating over a learned token space, where the distinction between automaton, grammar, & memory storage/retrieval is blurred. Rather than explicitly storing rules, the system implements them implicitly through its architecture, enabling the replay and recombination of stored sequences without relying on explicit symbolic guidance.
Critically, the model acquires long-term memory not through fixed architecture but through a dynamic inference-driven learning process we call replay. This mechanism enables the system to regenerate and reorganize previously learned structures, binding them into higher-level embedding neurons. These embeddings act as a sparse key-value memory (Gershman et al., 2025; Ba et al., 2016; Tyulmankov et al., 2021), supporting pattern completion and selective recall. Once replay is used to compress and store learned patterns, the lower-layer short-term memory can be forgotten or reused — a form of memory indirection that enables continual learning and concept recombination. In this way, the model achieves Turing-like memory dynamics (Graves et al., 2014, 2016): hierarchical information can persist even as local memory is overwritten, supporting compositional generalization and long-range inference.
A central strength of this framework lies in its interpretability. Each step in the learning process — from Hebbian updates to hierarchical token composition — is locally defined and analytically tractable. The structure and evolution of the model can be explicitly visualized through small toy examples, enabling direct inspection of how words are learned, compressed, and replayed. Unlike black-box models, the memory tensors and projector maps in each layer can be examined and manipulated, allowing one to trace the flow of information across layers and interpret model decisions in terms of compositional features. This makes the framework not only an effective generative model but also a platform for theoretical analysis and conceptual clarity — a rare intersection of symbolic transparency and neural dynamics.
Because of this, we have organized the first half of the paper (Sec 2-6) as both the model introduction & practical tutorial. Sec 2-3 discuss short-term memory (STM). Sec 4-5 discusses long-term memory (LTM) & their emergent key-value mechanisms. In Sec 7, we demonstrate how compositional STM can be creative. We touch on the deeper relationship between inductive biases and natural language structure in Sec 7.1. We end with an extended discussion.
@@ 2.1. Local Learning Events
where $v_{k}$ is the free token to be predicted, and $\tilde{v}^{(n-1)}_{\mu_{n-1}}$ is the learned projection of the $n{-}1$ trailing context tokens. Each such event represents a local time-step in the dynamics of the model, and triggers a Hebbian memory update:
@@ 2.2. Hierarchical Retokenization
To organize these events across scales, the model constructs a recursive sequence of projection maps $P_{n}$ that define higher-order features. Each projection promotes a pair $(\tilde{v}^{(n-1)},v)$ to a new token $\tilde{v}^{(n)}$, if it satisfies the condition:
@@
These retokenized features form the model’s internal vocabulary. Importantly, the projection maps are constrained by a smoothness condition: a new $n$-gram token is allowed only if all its subcomponents are already valid $(n{-}1)$-grams. This ensures that every token in the hierarchy is recursively decomposable — i.e retokenizable.
@@ 3. Retokenization and Emergent $n$-Point Structure
the recursive definition of projected tokens induces an implicit $n$-point interaction. Each projection is defined via Eqn 2 and applying this recursively from the basis layer ($n=1$) yields:
@@ 3.2. Left-Inference via Retokenization
To perform left-inference—predicting the initial token of a word given trailing context—we retokenize the energy function into a left-gauge form. For each layer $n$, the hierarchical Hopfield model learns 2-point correlations of the form:
@@ 3.3. Example: Learning from dogs, cats, and frogs
To illustrate how the hierarchical Hopfield model builds structure from scratch, we walk through the learning process for a minimal training set consisting of three words:
@@
We assume the model begins with a basis token set of dimension $d=26$, one for each letter of the English alphabet: $v_{j}\in\{v_{\bf a},\dots,v_{\bf z}\}$. The model learns from co-occurrence patterns in the training data, using local Hebbian updates following Eqn 1.
@@
$P_{3}^{\mu_{3},\mu_{2},j_{3}}=1\quad\text{for learned combinations}\quad\tilde{v}^{(3)}_{\mu_{3}}=\tilde{v}^{(2)}_{\mu_{2}}\,v_{j_{3}}.$
@@
(13) $P_{2}^{\mu_{2},j_{1},j_{2}}$ $=1\quad\text{for }(\mu_{2},j_{1}j_{2})\in\scriptstyle{\{({\bf do},{\bf d},{\bf o}),\ ({\bf og},{\bf o},{\bf g}),\ ({\bf gs},{\bf g},{\bf s}),\ ({\bf ca},{\bf c},{\bf a}),}$
@@
$P_{3}^{\mu_{3},\mu_{2},j_{3}}$ $=1\quad\text{for }(\mu_{3},\mu_{2},j_{3})\in\scriptstyle{\{({\bf dog},{\bf do},{\bf g}),\ ({\bf ogs},{\bf og},{\bf s}),\ ({\bf cat},{\bf ca},{\bf t}),}$
$\scriptstyle{\hskip 59.50012pt({\bf ats},{\bf at},{\bf s}),\ ({\bf fro},{\bf fr},{\bf o}),\ ({\bf rog},{\bf ro},{\bf g})\}}$
$P_{4}^{\mu_{4},\mu_{3},j_{4}}$ $=1\quad\text{for }(\mu_{4},\mu_{3},j_{4})\in\scriptstyle{\{({\bf dogs},{\bf dog},{\bf s}),\ ({\bf cats},{\bf cat},{\bf s}),\ ({\bf frog},{\bf fro},{\bf g})\}}$
Or equivalently written as a superposition of tensor product states (a.k.a an entangled state (Schollwöck, 2011)):

### v2

# Self-Organizing Language

@@ Abstract.
We introduce a novel paradigm of emergent local memory. It is a continuous-learning completely-parallel content-addressable memory encoding global order. It demonstrates how local constraints on uncoordinated learning can produce topologically protected memories realizing emergent symbolic order. It is therefore a neuro-symbolic bridge.
It further has the ability to produce human language without data, by exploiting its own self-organizing dynamics. It teaches us that words arise as a side-effect of emergent symbolic order, and that human language patterns at all structural levels reflect a universal mechanism of word formation (which is subregular). This work answers essential questions about the existence & origin of all the human language data.
Exotic AI, Hebbian, Physics-inspired Neural Nets, Interpretable Models, Hierarchical Compression, Language Emergence, Neuro-symbolic, Neuromorphic, Topological Protected Memory
@@ 1. Introduction
Language did not exist in the early universe, and yet it exists today. Thus learning language does not require language data. We demonstrate this here by constructing a model which converts noise into language.
By language, we refer to human language, as in the language which children learn and are responsible for. Specifically, we pre-assume a primative token set (such as an alphabet), and work entirely in the regime of human language patterns. [^1] The emergence of discrete categories from estimating continuous information has been studied by other authors (Boersma2020_NeuralNets4Phonology). These patterns, while seemingly arbitrary, have universal statistical structure quantifiable in terms of the distribution of these patterns: hierarchical, with finite word length, and a scale-invariant sub-word structure related to its readabilitywithoutspaces. This coincides with a multi-scale symbolic structure, a limit of accepted & well-formed sequences.
We thus define a valid human language model by its relationship to universals. A rigorous model must both (i) be grounded in physical reality, i.e local neural learning, and (ii) generate the universal structures of human language without pre-fitting them. Symbolic theories fail the first criterion: they describe high-level rules but are not realizations of local learning. Modern statistical language models (e.g LLM’s) fail both criteria: they are non-local pattern recognizers which fit pre-existing data but cannot account for why the data exists. What’s needed is a model which satisfies both.
The model we present satisfies these criteria. By model, we do not mean a single function, composed of other functions, whose parameters are optimized to minimize a global objective. Instead we refer to a set of locally coordinated rules, akin to a physical system, which continuously grows a model neuron-by-neuron. This places it in the same (local & objective-free) paradigm of model as Conway’s Game of Life (PaulRendell2011_ConwayTuringComplete; PaulRendell2014thesis_ConwayTuringComplete) or Lenia (Chan2019_Lenia), except in the domain of language and neural learning.
Despite being a highly uncoordinated learner, the system can represent global order. By global order we mean not just simple patterns, like ferromagnetic or antiferromagnetic alignment, but order rich enough to encode arbitrary strings—for instance, spins arranging to spell out the names of all the US presidents in sequence. In this sense, language itself is global order. Such order is built locally, through a hierarchical memory we call the retokenizer. Even when trained on noise, the retokenizer self-organizes into a structure that encodes strings with universal sub-word distributions and finite word length. These strings act as symbols: terminal nodes of a bifix-free [^2] Preffix & suffix free, meaning that cat & catch cannot both distinguished in the same instance of memory. Only the longer one is retrieved. directed acyclic graph (DAG), whose topology guarantees their retrieval. In this way, symbolic behavior is not imposed but emerges from the topological character of the memory.
[…]
$a_{\bf you}=\tilde{v}^{(3)}_{\bf you}\oplus\big(\tilde{v}^{(2)}_{\bf yo}+\tilde{v}^{(2)}_{\bf ou}\big)\oplus\big(v_{\bf y}+v_{\bf o}+v_{\bf u}\big).$
The formation of object tokens marks the transition from short-term correlations to long-term symbolic memory. Each object token is completely independent, storing the feature tokens of the retokenizer (whose memory can be forgotten). Inference (next-token prediction) now follows a key-value relationship, an emergent attention mechanism with knowledge-graph behavior akin to transformer embeddings. Crucially, object tokens can themselves be correlated, producing higher-order object tokens. Thus the architecture is not fixed: new memories can be formed without erasing old ones, and the computational class itself evolves with continued learning.
The complete model is not a singular thing, but a non-static continuous-learning completely-parallel content addressable memory encoding global order. Therefore, what we present is a novel paradigm of emergent local memory. Such a model is essential for understanding how the universal constraints of the brain are themselves responsible for the observed macroscopic language order. Thus it is also an effective human language model, and the only one we are aware of which meets the minimal criteria laid out in this intro.
In this paper, we focus on the low-data regime, which is the essential regime for answering key questions related to human memory. Conveniently, local models are analytically tractable & thus highly interpretable. It’s parallel nature makes it so that the pedagogical toy models we present are actually representative of how the model can learn in practice – bit by bit.
The combination of analytic simplicity and multi-scale symbolic structure makes it possible to construct model initial states with syntax-like inter-categorical relationships and hierarchy. As a demonstration, we provide free code [^4] https://gitlab.com/emla-group/ababot where such initial states can be specified directly, allowing symbolic theories of English (or other languages) to be hand-coded and then grown through local learning.
This manuscript is organized into two main parts: The first part encompasses the tensor machinery of short-term memory (STM), which builds to a discussion on the structural origins of human subword structure (Sec 4). The second half covers the mechanisms responsible for forming long-term memory (LTM)—starting first with a summary (Sec 5), followed by mathematical details. We simplify the math & discuss the high-level organization of the code in Sec 7. Then discuss Syntax in Sec 8. We end with some toy examples in Sec 9.
@@ 2. Model Summary: Local Events and Hierarchical Retokenization
2.1. Preliminary
[…]
We further define a token pool, which is an ordered list of tokens used as a basis, e.g $v_{\bf a},v_{\bf b}]$. This allows us to construct a dense representation, while keeping track of which vector rows correspond to which tokens. For our dimension $d=2$ example pool:
$v_{\bf a}$ $=$ $\begin{pmatrix}1&0\end{pmatrix}$
$v_{\bf b}$ $=$ $\begin{pmatrix}0&1\end{pmatrix}$,
and use Latin indices ($j,k,l\in\{{\bf a},{\bf b}\}$) to write $v_{k}\in\{v_{\bf a},v_{\bf b}\}$. Tensor products of tokens (written $v_{\bf a}\otimes v_{\bf b}\equiv v_{\bf a}v_{\bf b}$) themselves form their own $d^{2}$ dimensional orthonormal set,
$v_{\bf a}v_{\bf a}$ $=$ $\begin{pmatrix}1&0&0&0\end{pmatrix}$
$v_{\bf a}v_{\bf b}$ $=$ $\begin{pmatrix}0&1&0&0\end{pmatrix}$
$v_{\bf b}v_{\bf a}$ $=$ $\begin{pmatrix}0&0&1&0\end{pmatrix}$
$v_{\bf b}v_{\bf b}$ $=$ $\begin{pmatrix}0&0&0&1\end{pmatrix}$.
One could even define $v_{\bf ab}^{(2)}\equiv v_{\bf a}v_{\bf b}$ to denote that they are higher dimensional vectors. Higher-order tensor products form the generic basis for representing tensors, e.g
(1) $T^{(n)}=\sum_{l,\cdots,j,k}T_{l,\cdots,j,k}^{(n)}v_{l}\otimes\cdots\otimes v_{j}\otimes v_{k},$
where $T_{l,\cdots,j,k}^{(n)}$ is the matrix element. A generic tensor $T$ is a direct sum (concatenation) over all subspaces,
(2) $T=T_{(1)}\oplus T_{(2)}\oplus\cdots\oplus T_{(n)}\oplus\cdots,$
equivalently
(3) $=\sum_{n}\sum_{l,\cdots,j,k}T_{l,\cdots,j,k}^{(n)}v_{l}\otimes\cdots\otimes v_{j}\otimes v_{k}.$
Note that context refers to a list of tokens, for either learning or autoregressively predicting subsequent tokens. A generic context $v(1),v(2),\cdots,v(n)]$ can be mapped to a direct sum over tensor products
(4) $T_{\text{context}}=$
$v(n)\oplus\big[v(n-1)v(n)\big]\oplus\cdots\oplus\big[v(1)v(2)\cdots v(n)\big]$
where subspaces beyond length $n$ are implicitly padded by zeros. Tensors act on context via a dot product, i.e a similarity check. For example, $T_{\text{context}}=v_{\bf a}\oplus v_{\bf b}v_{\bf a}\oplus v_{\bf a}v_{\bf b}v_{\bf a}$, reduces to the sum
(5) $T\cdot T_{\text{context}}=T_{\bf a}^{(1)}+T_{\bf ba}^{(2)}+T_{\bf aba}^{(3)}.$
Note that the quantity $n$ is the coordination length. Tensor products of length $n$ require all $n$ tokens to be non-zero in order for the whole product to be finite—hence they can be understood as encoding conditional AND(&) statements.
In general, the dimension of a tensor product grows exponentially in the number of products (nn). [^5] One can understand $T$ as living in a real-valued Fock space. Note that we can reduce the memory overhead by truncating the exponentially-large tensor product space, defining a reduced basis set written $\tilde{v}_{\mu_{n}}^{(n)}$, where $d_{n}\equiv\text{dim}(\mu_{n})\ll d^{n}$. For example ($d=2$), keeping ab & ba:
$\tilde{v}_{\bf ab}^{(2)}$ $=$ $\begin{pmatrix}1&0\end{pmatrix}$
$\tilde{v}_{\bf ba}^{(2)}$ $=$ $\begin{pmatrix}0&1\end{pmatrix}$.
We will show this effectuates a low-rank compression,
(6) $\text{compressed}\big(T^{(n)}\big)=\sum_{\mu_{n-1},k}T_{\mu_{n-1},k}^{(n)}\tilde{v}^{(n-1)}_{\mu_{n-1}}\otimes v_{k},$
where the trailing token products are merged into a single compressed vector. This reduces the tensor product to effectively $2$-point coordination (i.e locally coordinated).
The sections which follow demonstrates how to grow low-rank tensors from local learning events, which directly map to $n$-points in the uncompressed $d^{n}$ space.
Throughout this paper we use standard tensor notation. Bold font distinguish labels (a,b,c) from free indices ($k,\mu_{n},\alpha$). Some indices ($\mu_{n}$) carry sub-indices ($n$) indicating layer.
@@ 2.2. Local Learning Events
where $v_{k}$ is the free token (a vector of dimension $d$) to be predicted, and $\tilde{v}^{(n-1)}_{\mu_{n-1}}$ is the learned projection of the n1n{-}1 trailing context tokens. Each such event represents a local time-step in the dynamics of the model, and triggers a Hebbian memory update:
@@ 2.3. Hierarchical Retokenization
To organize these events across scales, the model constructs a recursive sequence of projection maps $P_{n}$ that define higher-order features. Each projection promotes a pair $(\tilde{v}^{(n-1)},v)$ to a single token $\tilde{v}^{(n)}$, if it satisfies the condition:
@@
These retokenized features form the model’s internal vocabulary.
The $P_{n}$ are grown layer-by-layer over a finite window of training tokens. Growths form if $g_{\mu_{n-1},k}^{(n)}>\epsilon_{n}$ for cutoff $\epsilon_{n}$. Importantly, the projection maps are constrained by a smoothness condition: a new $n$-gram token is allowed only if all its subcomponents are already valid (n1)(n{-}1)-grams. This ensures that every token in the hierarchy is recursively decomposable — i.e retokenizable.
@@ 3. Retokenization and Emergent $n$-Point Structure
the recursive definition of projected tokens induces an implicit $n$-point interaction. Each projection is defined via Eqn 8 and applying this recursively from the basis layer ($n=1$) yields:
@@ 3.2. Left-Inference via Retokenization
To perform left-inference—predicting the initial token of a word given trailing context—we retokenize the energy function into a left-gauge form. For each layer $n$, the hierarchical Hebbian-Ising model learns 2-point correlations of the form:
@@ 3.3. Example: Learning from dogs, cats, and frogs
To illustrate how the hierarchical Ising model builds structure from scratch, we walk through the learning process for a minimal training set consisting of three words:
@@
We assume the model begins with a basis token set of dimension $d=26$, one for each letter of the English alphabet: $v_{j}\in\{v_{\bf a},\dots,v_{\bf z}\}$. The model learns from co-occurrence patterns in the training data, using local Hebbian updates following Eqn 7.
@@
$P_{3}^{\mu_{3},\mu_{2},j_{3}}=1\quad\text{for learned combinations}\quad\tilde{v}^{(3)}_{\mu_{3}}\cong\tilde{v}^{(2)}_{\mu_{2}}\,v_{j_{3}}.$
@@
(19) $P_{2}^{\mu_{2},j_{1},j_{2}}$ $=1\qquad\text{for }(\mu_{2},j_{1}j_{2})\in\scriptstyle{\{({\bf do},{\bf d},{\bf o}),\ ({\bf og},{\bf o},{\bf g}),\ ({\bf gs},{\bf g},{\bf s}),\ ({\bf ca},{\bf c},{\bf a}),}$
@@
$P_{3}^{\mu_{3},\mu_{2},j_{3}}$ $=1\qquad\text{for }(\mu_{3},\mu_{2},j_{3})\in\scriptstyle{\{({\bf dog},{\bf do},{\bf g}),\ ({\bf ogs},{\bf og},{\bf s}),\ ({\bf cat},{\bf ca},{\bf t}),}$
$\scriptstyle{\hskip 69.65302pt({\bf ats},{\bf at},{\bf s}),\ ({\bf fro},{\bf fr},{\bf o}),\ ({\bf rog},{\bf ro},{\bf g})\}}$
$P_{4}^{\mu_{4},\mu_{3},j_{4}}$ $=1\qquad\text{for }(\mu_{4},\mu_{3},j_{4})\in\scriptstyle{\{({\bf dogs},{\bf dog},{\bf s}),\ ({\bf cats},{\bf cat},{\bf s}),\ ({\bf frog},{\bf fro},{\bf g})\}}$
Or equivalently written as a superposition of tensor product states (a.k.a an entangled state (mps_Schollwock2011)):
@@
In formal language terms, the feature maps act as finite-state automata which accepts strings whose complexity is subregular [^6] Formally, regular languages are recognized by finite-state automata that lack explicit memory for counting or nesting (Hopcroft&Motwani&Ullman2001book_AutomataTheoryLanguages&Computation)..
Figure 1. Local interactions ($\tilde{v}^{(n-1)}\otimes v$) drive new model growths ($\tilde{v}^{(n)}$) when input events align with the allowed hierarchical DAG structure. Stored strings are induced global order parameters ($v\otimes v\otimes\cdots\otimes v\rightarrow\tilde{v}^{(n)}$), each living as a terminal node of the DAG. Each growth is compressed/decompressed as a matrix product state (MPS).
3.3.1. Topology of the Ground State Automaton